# TripNet: Learning Large-scale High-fidelity 3D Car Aerodynamics with Triplane Networks

### v1

@@ 4 Results
We evaluate our method on DrivAerNet (Elrefaie et al., 2024a) and DrivAerNet++ (Elrefaie et al., 2024d) datasets across aerodynamics tasks including drag coefficient predictions, surface field pressure, and 3D flow field simulations. These datasets were chosen because they are the largest industry-standard datasets available with extensive geometric variations and high-fidelity simualations, making them well-suited for data-driven methods. The drag model is trained on a single RTX 4090 GPU for 1 hour, while the surface and 3D flow field models are trained on four H100 GPUs over a period of 1.5 days.
@@ 4.1 Aerodynamic Drag Prediction
The superior performance of TripNets stems from the triplane representation, which preserves full geometric information in an implicit form, unlike traditional mesh, voxel, or point-cloud methods that suffer from downsampling losses. This richer and more expressive feature space enables precise aerodynamic predictions using a lightweight CNN model. While prior approaches attempted to overcome geometric limitations with handcrafted features, such as multi-view image rendering (Song et al., 2023), these methods introduce human bias and are constrained by predefined viewpoints. In contrast, TripNets autonomously learn task-specific representations, ensuring optimized and highly structured features for aerodynamic prediction, leading to enhanced accuracy and robustness.
Table 2: Performance comparison on surface pressure field prediction across various models. The evaluation is conducted on the unseen test set of both the DrivAerNet dataset, which includes 600 industry-standard car designs, and the DrivAerNet++ dataset, which includes 1200 industry-standard car designs.
@@
DrivAerNet RegDGCNN (Elrefaie et al., 2024a) 9.01 1.58 13.09 28.49 26.29
FigConvNet (Choy et al., 2025) 4.38 1.13 5.73 20.98 18.59
@@
DrivAerNet++ RegDGCNN (Elrefaie et al., 2024a) 8.29 1.61 10.81 27.72 26.21
FigConvNet (Choy et al., 2025) 4.99 1.22 6.55 20.86 21.12
@@
4.2 Surface Field
Figure 2: Comparison of pressure predictions and errors for car design N_S_WW_WM_025 (notchback with smooth underbody, with wheels, and with mirrors) from the unseen test set. The first row shows the input mesh followed by predictions from RegDGCNN, FigConvNet, TripNet (ours), and the ground-truth CFD. The second row highlights the absolute difference between the predictions and ground-truth CFD, visualizing the absolute error distribution for each model.
@@ 4.2 Surface Field
Figure 2 illustrates the pressure predictions and error distributions for the car design N_S_WW_WM_025, demonstrating the accuracy of different models, including our TripNet, in comparison to ground-truth CFD results. Similarly, Table 2 shows that TripNets outperform previous state-of-the-art models in predicting surface pressure fields on both the DrivAerNet and DrivAerNet++ datasets.
@@ 4.3 Volume Flow Field
Error (%$\downarrow$ Error (%$\downarrow$
@@
While triplane features are initially extracted to encode surface geometry information, they inherently capture global geometric context, including regions corresponding to empty space around the car. When revisiting the geometry reconstruction task used to fit the triplane, we observe that the triplane features not only represent surface properties but also provide valuable information about the spatial structure of the surrounding flow field. This insight allows us to extend triplane processing beyond the car surface, leveraging its information-rich representation to predict full 3D velocity fields near the car. Table 3 shows the performance metrics on volumetric flow field predictions observed between TripNet predictions and the simulation, averaged over the test set of DrivAerNet++. TripNets achieve a 7.15% $L_{1}$ error and 10.7% $L_{2}$ error for $U_{x}$ prediction. A higher $L_{1}$ and $L_{2}$ error is observed for $U_{y}$ and $U_{z}$ due to their sparse distribution and lower magnitude compared to $U_{x}$. The training, validation, and test dataset for volumetric flow field prediction consists of 8,000 simulations, amounting to approximately 16TB of high-fidelity CFD data.
Figure 3 presents a comparison of 3D flow field predictions for velocity magnitude ($U$) in three cross-sectional planes of the car design E_S_WWC_WM_094. The figure visualizes the ground truth CFD results, our model’s predictions, and the corresponding absolute error across the symmetry plane ($y=0$), the wake region ($x=4$), and a vertical slice cutting the car body ($z=1$). Our model accurately predicts the wake flow in the rear of the car as well as the wake generated by the side mirrors, demonstrating its effectiveness in capturing complex aerodynamic flow structures.
@@ 5 Conclusion
Our approach efficiently encodes aerodynamic properties into triplane features, enabling multiple predictive tasks. For drag coefficient prediction, we utilize compressed triplane features obtained through channel-wise max, min, and average pooling, feeding them into a lightweight CNN-based model. On the DrivAerNet dataset, our model achieves an $R^{2}$ score of 0.97, requiring only one hour of training and 0.01 seconds for inference per car. For surface pressure and WSS estimation, we employ raw triplane features, processed through a U-Net, allowing field values to be queried at specific spatial coordinates.
@@
Appendix A Triplane Fitting
@@
We followed the NDF paper to fit the triplane and adopted a two-stage training to fit the triplanes for cars in the DrivAerNet++ dataset. In Stage 1, we randomly picked 500 cars and jointly trained the triplane of each car and the shared MLP decoder. We train the first stage with 500 epochs and a learning rate of 1e-3. In Stage 2, the MLP decoder is shared and we fit the triplane of each object individually across multiple machines for the sake of efficiency. The fit of each takes  20 seconds to converge with 70 epochs. The triplane obtained from geometry reconstruction task and used for aerodynamics tasks is shown in Fig. A1.
@@
Figure A1: Visulization of the raw triplane features of F_D_WM_WW_1282 (fastback with detailed underbody, with wheels, and with mirrors) used for surface field and 3D flow field predictions.
Appendix B Triplane Resolution
Figure A2: 3D Geometry reconstructed by triplanes with different dimensions.
@@ Appendix B Triplane Resolution
To evaluate the impact of triplane resolution on 3D geometry reconstruction, we compare different triplane resolutions and analyze their effects on reconstruction quality. Figure. A2 illustrates the reconstructed 3D geometries at various triplane resolutions, while Table. A1 provides quantitative metrics, including precision, recall, and F1-score, to assess reconstruction fidelity. Our results show that increasing triplane resolution improves reconstruction accuracy. At a lower resolution of $32\times 64\times 64$, the reconstructed geometry retains the overall shape but lacks fine surface details. As the resolution increases to $32\times 128\times 128$, the details become more refined, and the precision, recall, and F1-score improve. At $32\times 256\times 256$, the reconstruction achieves near-perfect fidelity, closely matching the original mesh, with an F1-score of 0.9936.
@@
However, higher triplane resolutions come with increased memory and computational costs. A resolution of $32\times 128\times 128$ offers a good balance between reconstruction accuracy and efficiency. These results highlight the trade-off between resolution and computational efficiency, which should be carefully considered when selecting the appropriate triplane representation for downstream tasks.
@@
Table A1: Impact of triplane resolution on geometry reconstruction quality. Higher triplane resolutions lead to improved precision, recall, and F1 score, indicating more accurate and detailed geometry representation.
@@
Table A2 presents the performance metrics of our Triplane network across different resolutions for the surface pressure field task on the DrivAerNet dataset. As the triplane resolution increases from $32\times 64\times 64$ to $32\times 128\times 128$, there is a notable improvement in performance metrics, including Mean Squared Error (MSE), Mean Absolute Error (MAE), Maximum Absolute Error, Relative L2 Error, and Relative L1 Error. However, further increasing the resolution to $32\times 256\times 256$ does not yield additional performance gains, indicating a potential saturation point in the network’s ability to leverage higher-resolution inputs for this specific task.
Table A2: Performance of the Triplane network using different triplane resolutions on surface pressure field task on DrivAerNet dataset.
@@
Appendix C Additional Results
C.1 Wall Shear Stress Field
Table A3 presents the performance comparison of different models on surface wall shear stress (WSS) field prediction, evaluated on the unseen test set of the DrivAerNet++ dataset. The mean squared error (MSE) and mean absolute error (MAE) are computed over all three components ($x,y,$ and $z$) to assess the overall accuracy of the predicted WSS vectors. TripNet outperforms previous state-of-the-art methods, achieving the lowest mean absolute error (2.15 ×$\times 10^{-1}$) and relative $L_{1}$ error (17.18%). Compared to RegDGCNN and FigConvNet, TripNet consistently achieves better accuracy, demonstrating its effectiveness in learning fine-grained WSS distributions. In addition, Figures A5 and A6 provide a detailed comparison of WSS predictions and errors across multiple deep learning models for three car designs selected from the unseen test set.
Table A3: Performance comparison on surface wall shear stress (WSS) field prediction across various models. The evaluation is conducted on the unseen test set of the DrivAerNet++ dataset, which includes 1200 industry-standard car designs.

### v2

@@ 2.3 Surface Field and Volume Flow Field Tasks
2.4 Specialization Vs Generalization
In this paper, we evaluate and compare model performance on both DrivAerNet and DrivAerNet++. DrivAerNet serves as a specialization benchmark, focusing on a single car category with limited geometric variation, making it suitable for studying performance in constrained design spaces and class-specific predictions. In contrast, DrivAerNet++ presents a significantly more challenging generalization task, encompassing multiple car categories and a broader diversity of designs. This enables us to assess each model’s ability to generalize across varying shapes and aerodynamic configurations, transitioning from class-specific to cross-class predictions.
@@ 3 Results
We evaluate our method on DrivAerNet Elrefaie et al. (2024b) and DrivAerNet++ Elrefaie et al. (2024d) datasets across aerodynamics tasks including drag coefficient predictions, surface field pressure, and 3D flow field simulations. These datasets were chosen because they are the largest industry-standard datasets available with extensive geometric variations and high-fidelity simualations, making them well-suited for data-driven methods. The drag model is trained on a single RTX 4090 GPU for 1 hour, while the surface and 3D flow field models are trained on four H100 GPUs over a period of 1.5 days.
@@ 3.1 Aerodynamic Drag Prediction
The superior performance of TripNets stems from the triplane representation, which preserves full geometric information in an implicit form, unlike traditional mesh, voxel, or point-cloud methods that suffer from downsampling losses. This richer and more expressive feature space enables precise aerodynamic predictions using a lightweight CNN model. While prior approaches attempted to overcome geometric limitations with handcrafted features, such as multi-view image rendering Song et al. (2023), these methods introduce human bias and are constrained by predefined viewpoints. In contrast, TripNets autonomously learn task-specific representations, ensuring optimized and highly structured features for aerodynamic prediction, leading to enhanced accuracy and robustness.
Table 2: Performance comparison on drag coefficient prediction across various models using different data representations, including point clouds, graphs, and our triplane representation. The evaluation is conducted on the DrivAerNet test set, which includes 600 industry-standard car designs. Models are ranked based on $R^{2}$ performance.
Model MSE $\downarrow$ MAE $\downarrow$ Max AE $\downarrow$ $R^{2}$uparrow
(×$(\times 10^{-5})$ (×$(\times 10^{-3})$ (×$(\times 10^{-2})$
PointNet Qi et al. (2017a) 12.0 8.85 10.18 0.826
GCNN Kipf and Welling (2016) 10.7 7.17 10.97 0.874
PointNet++ Qi et al. (2017b) 7.813 6.755 3.463 0.896
RegDGCNN Elrefaie et al. (2024b) 8.01 6.91 8.80 0.901
PointBERT Yu et al. (2022) 6.334 6.204 2.767 0.915
DeepGCN Li et al. (2019) 6.297 6.091 3.070 0.916
MeshGraphNet Pfaff et al. (2020) 6.0 6.08 2.965 0.917
AssaNet Qian et al. (2021) 5.433 5.81 2.39 0.927
PointNeXt Qian et al. (2022) 4.577 5.2 2.41 0.939
FIGConvNet Choy et al. (2025) 3.225 4.423 2.134 0.957
TripNet (Ours) 2.602 4.030 1.268 0.972
3.2 Surface Field
Table 3: Performance comparison on surface pressure field prediction across various models. The evaluation is conducted on the unseen test set of both the DrivAerNet and DrivAerNet++ dataset, which includes 600 and 1200 industry-standard car designs
@@ 3.2 Surface Field
(×$(\times 10^{-2})$ (×$(\times 10^{-1})$ (%) $\downarrow$ (%$\downarrow$
DrivAerNet RegDGCNN Elrefaie et al. (2024b) 9.01 1.58 13.09 28.49 26.29
FigConvNet Choy et al. (2025) 4.38 1.13 5.73 20.98 18.59
Transolver Wu et al. (2024a) 5.37 1.28 5.89 22.52 21.31
@@
DrivAerNet++ RegDGCNN Elrefaie et al. (2024b) 8.29 1.61 10.81 27.72 26.21
FigConvNet Choy et al. (2025) 4.99 1.22 6.55 20.86 21.12
Transolver Wu et al. (2024a) 7.15 1.41 7.12 23.87 22.57
@@
Table 4: Performance comparison on surface wall shear stress field prediction across various models on the DrivAerNet and DrivAerNet++ dataset
[…]
(×$(\times 10^{-2})$ (×$(\times 10^{-1})$ (%) $\downarrow$ (%$\downarrow$
DrivAerNet RegDGCNN Elrefaie et al. (2024b) 12.13 1.94 11.53 34.70 35.53
FigConvNet Choy et al. (2025) 9.10 1.79 7.85 22.01 18.59
Transolver Wu et al. (2024a) 8.59 1.69 8.10 21.51 16.92
TripNet (Ours) 8.14 1.66 7.47 21.80 17.39
DrivAerNet++ RegDGCNN Elrefaie et al. (2024b) 13.82 3.64 11.54 36.42 36.57
FigConvNet Choy et al. (2025) 9.86 2.22 7.84 22.32 17.65
Transolver Wu et al. (2024a) 8.95 2.06 8.45 22.49 17.65
TripNet (Ours) 9.52 2.15 8.14 22.07 17.18
Figure 2: Comparison of pressure predictions and errors for car design E_S_WWC_WM_094 (estateback with smooth underbody, wheels closed, and with mirrors) from the unseen test set. The first row shows the input mesh followed by predictions from RegDGCNN, FigConvNet, Transolver, TripNet (ours), and the ground-truth CFD. The second row highlights the absolute difference between the predictions and ground-truth CFD, visualizing the absolute error distribution for each model.
@@
Figure 2 illustrates the pressure predictions and error distributions for the car design E_S_WWC_WM_094 from the unseen test set, demonstrating the accuracy of different models, including our TripNet, in comparison to ground-truth CFD results. TripNet demonstrates state-of-the-art performance in both wall shear stress and surface pressure prediction tasks across the DrivAerNet and DrivAerNet++ datasets. As shown in Table 3 and Table 4, TripNet achieves the lowest or second-lowest values across all key evaluation metrics—including MSE, MAE, Max AE, $\ell_{1}$ and $\ell_{2}$ error compared to established baseline models such as RegDGCNN, FigConvNet, and Transolver. These results indicate that TripNet matches or exceeds the performance of current state-of-the-art approaches, highlighting its effectiveness in high-fidelity aerodynamic field prediction tasks.
@@ 3.3 Volume Flow Field
(%$\downarrow$ (%$\downarrow$
@@
While triplane features are initially extracted to encode surface geometry information, they inherently capture global geometric context, including regions corresponding to empty space around the car. When revisiting the geometry reconstruction task used to fit the triplane, we observe that the triplane features not only represent surface properties but also provide valuable information about the spatial structure of the surrounding flow field. This insight allows us to extend triplane processing beyond the car surface, leveraging its information-rich representation to predict full 3D velocity fields near the car. Table 5 shows the performance metrics on volumetric flow field predictions observed between TripNet predictions and the simulation, averaged over the test set of DrivAerNet++. TripNets achieve an $\ell_{1}$ error of 7.15% and $\ell_{2}$ error of 10.7% for $U_{x}$ prediction. A higher $\ell_{1}$ and $\ell_{2}$ error is observed for $U_{y}$ and $U_{z}$ due to their sparse distribution and lower magnitude compared to $U_{x}$. The training, validation, and test dataset for volumetric flow field prediction consists of 8,000 simulations, amounting to approximately 16TB of high-fidelity CFD data.
Figure 3 presents a comparison of 3D flow field predictions for velocity magnitude ($U$) in three cross-sectional planes of the car design E_S_WWC_WM_094. The figure visualizes the ground truth CFD results, our model’s predictions, and the corresponding absolute error across the symmetry plane ($y=0$), the wake region ($x=4$), and a vertical slice cutting the car body ($z=1$). Our model accurately predicts the wake flow in the rear of the car as well as the wake generated by the side mirrors, demonstrating its effectiveness in capturing complex aerodynamic flow structures. Additional comparison between TripNet and CFD results is provided in Appendix.  C.4.
Figure 4: Scalability comparison of TripNet, FigConvNet, and Transolver across different point resolutions. (a) Relative L2 Error vs. Number of Points. (b) GPU Memory Usage vs. Number of Points; (c) Inference Time vs. Number of Points
@@ 4 Conclusion
Our approach efficiently encodes aerodynamic properties into triplane features, enabling multiple predictive tasks. For drag coefficient prediction, we utilize compressed triplane features obtained through channel-wise max, min, and average pooling, feeding them into a lightweight CNN-based model. On the DrivAerNet dataset, our model achieves an $R^{2}$ score of 0.97, requiring only one hour of training and 0.01 seconds for inference per car.
TripNet outperforms or matches state-of-the-art models on both pressure and wall shear stress prediction tasks across the DrivAerNet and DrivAerNet++ datasets. For pressure prediction, TripNet achieves the lowest relative $\ell_{2}$ errors of $20.35\%$ (DrivAerNet) and $20.05\%$ (DrivAerNet++), compared to $20.98\%$ and $20.86\%$ for FigConvNet, and $22.52\%$ and $23.87\%$ for Transolver. On wall shear stress prediciton, TripNet also leads with relative $\ell_{2}$ errors of 22.07%, while FigConvNet reports 22.32%, and Transolver lags behind at 22.49% on DrivAerNet++ dataset. As shown in Figure  4, TripNet offers significantly improved computational efficiency and scalability: TripNet uses 4.1$\times$ less GPU memory and is 1.4$\times$ faster in inference compared to FigConvNet at the highest resolution (131,072 points). Unlike FigConvNet and Transolver, whose GPU memory usage grows with input resolution, TripNet maintains nearly constant memory usage and inference time owing to its resolution-independent triplane representation. In addition, for volumetric flow field prediction, the structured and dense triplane feature representation used by TripNet is inherently more suitable than the sparse point cloud representation employed by FigConvNet, enabling more accurate and complete predictions across the full 3D domain.
@@
5 Limitation and Future Work
Despite its strong performance, our method has several limitations. The triplane fitting is performed separately for each geometry, so the model’s ability to generalize to topologically distinct shapes (e.g., F1 cars) or different flow regimes beyond the training Reynolds number remains unverified. Additionally, the model does not explicitly enforce physical constraints such as mass conservation or turbulence-model consistency, which may lead to non-physical artifacts, particularly in regions with sparse training data. Future work will explore incorporating physics-informed constraints and training on more diverse geometries and flow conditions to improve generalization and ensure physically consistent predictions.
Appendix A Related Work
RANS for Initial Design Stages
Reynolds-Averaged Navier-Stokes (RANS) simulations are a fundamental approach in industrial fluid dynamics, particularly valuable in the early and conceptual design stages due to their balance of accuracy and computational efficiency. Recent datasets for aerodynamic analysis have employed RANS simulations, showcasing a wide range of dataset sizes and mesh resolutions. These include 500 designs in Usama et al. Usama et al. [2021], 889 in Umetani et al. Umetani and Bickel [2018], 1,000 in Gunpinar et al. Gunpinar et al. [2019], and 1,400 in Remelli et al. Remelli et al. [2020]. Larger datasets include 2,474 designs in Song et al. Song et al. [2023] and 1,121 in Trinh et al. Trinh et al. [2024], with mesh resolutions ranging from 130k in Gunpinar et al. Gunpinar et al. [2019] to 3M in Song et al. Song et al. [2023]. The most recent DrivAerNet Elrefaie et al. [2024b] and DrivAerNet++ Elrefaie et al. [2024d] datasets were published to address the trade-off between large dataset sizes required for training deep learning models and the simulation fidelity necessary for industrial applications. These datasets comprise 8,000 designs with high-resolution meshes, providing an unparalleled level of detail for evaluating machine learning-based CFD approaches. This combination ensures robust performance benchmarking and facilitates the development of more accurate predictive models.
Deep Learning Models for Simulation:
Song et al Song et al. [2023] adopts a new 3D representation using stacked depth and normal renderings. Six single views, front, rear, top, bottom, right and left, are generated and integrated into a single image to capture the geometric features comprehensively. CNN-based and transformer-based models are used to learn the features from the 2D representation and predict drag coefficients. Jacob et al. Jacob et al. [2021] introduced a method using Signed Distance Fields (SDF) in combination with a U-Net archicture to simultaneously predict both the drag and velocity fields. This representation captures surface proximity and enables spatially-aware predictions across the flow field.
Graph Neural Networks (GNNs) have emerged as one of the most popular approaches for learning-based CFD simulations due to their ability to handle irregular geometries and unstructured meshes. However, several challenges remain in their application to large-scale, high-fidelity aerodynamic datasets. MeshGraphNets (MGNs) have been extensively studied for their generalization capabilities in fluid dynamics, particularly for unseen geometries Schmöcker et al. [2024]. A comparative analysis of computational performance on both CPU and GPU systems demonstrated a 6.5x speedup in inference on a CPU and a 100.8x speedup on a GPU, significantly reducing training and inference times for replicating full CFD simulations. However, this study primarily focuses on fluid simulations for standard geometries, utilizing meshes with $\mathcal{O}(10^{3})$ nodes. Two key limitations of this approach are (1) the use of relatively simple geometries (e.g., cylinders) and (2) the restricted number of nodes, again at $\mathcal{O}(10^{3})$. To address these scalability limitations, X-MeshGraphNet Nabian et al. [2024] presents a multi-scale, mesh-free extension of MGNs. It partitions large graphs and incorporates halo regions for inter-partition message passing. This design supports distributed training with equivalent full-graph performance. X-MeshGraphNet also constructs graphs directly from STL geometry files by generating surface or volumetric point clouds, connected via k-nearest neighbors. It further builds multi-resolution graphs to support long-range interactions, enabling real-time simulation without requiring mesh generation at inference. Despite the complexity, it maintains the predictive accuracy of full-graph GNNs.
Transformer-based solvers such as Transolver Wu et al. [2024a] enable PDE resolution on general geometries, operating on meshes from $2^{10}$ to $2^{15}$ nodes. Although they outperform linear transformer variants, scaling to industrial-scale datasets like DrivAerNet++, which contains millions of nodes, remains a challenge.
Neural operator is a family of machine learning models designed the approximate the solutions of PDEs. Unlike neural networks that map inputs to outputs, neural operators are trained to approximate the mapping between function spaces. Among the popular architectures, the Fourier Neural Operator (FNO)  Li et al. [2020] utilizes the Fourier transform to represent the kernel integration and solves PDEs in the spectral domain, achieving computational efficiency on regular grids by leveraging Fast Fourier Transforms (FFT). The Graph Neural Operator (GNO)  Li et al. [2024] adopts a graph representation for kernel integration, making it highly suitable for handling complex geometries and irregular domains.
To summarize the capabilities of these models, Table. 1 in the main paper provides a comparative overview of representative 3D deep learning architectures for solving Navier-Stokes equations. The table outlines geometry representations, prediction fashion, and scalability features. TripNet distinguishes itself by using a triplane-based implicit representation that supports query-based predictions beyond the input points and maintains memory efficiency across resolutions.
Implicit and Neural Representations:
Implicit and neural representations Serrano et al. [2023], Catalani et al. [2024] offer a continuous alternative to traditional discrete approaches in CFD simulations. Unlike GNNs and point cloud-based models, which provide solutions only at discrete mesh nodes, implicit representations enable querying the solution at any point in space (and time, if included as an input feature). This flexibility allows for higher-resolution predictions without being constrained by the mesh structure, making them particularly advantageous for modeling complex aerodynamic flows. de Vito et al. de Vito et al. [2024] presents an implicit neural representation framework for accurate 3D CFD flow field prediction, leveraging a coordinate-based MLP (backbone-net) to provide a discretization-agnostic, infinite-resolution representation of flow fields for domains with arbitrary topology. Combined with a hyper-net that maps blade surface geometry to the backbone-net’s parameters, this approach offers a memory-efficient, highly accurate proxy to CFD solvers, demonstrating strong potential for industrial applications like wing and turbine blade design.
Appendix B Models Description
The TripNet is trained on 4 H100 GPUs with 200 epochs. We use a batch size of 32, and for each surface field predictions, we sample 200,000 points in the training. All the other baseline models are trained on one A100 GPU. For a fair comparison, we also trained TripNet on one A100 GPU with a batch size of 8 and the 200,000 sampled points. The training setup is summarized in Table  1.
Table 1: Training setup of TripNet and baseline models on one A100 GPU
Model parameters Epochs Optimizer Batch size learning rate
RegDGCNN 5,750,820 150 Adam 12 1e-3
FigConvNet 77,649,570 200 Adam 1 1e-3
Transolver 3,860,033 200 Adam 1 1e-3
TripNet 24,122,672 200 Adam 32 1e-3
B.1 TripNet Overview
B.1.1 Triplane Fitting
@@
We followed the NDF paper to fit the triplane and adopted a two-stage training to fit the triplanes for cars in the DrivAerNet++ dataset. In Stage 1, we randomly picked 500 cars and jointly trained the triplane of each car and the shared MLP decoder. We train the first stage with 500 epochs and a learning rate of 1e-3. In Stage 2, the MLP decoder is shared and we fit the triplane of each object individually across multiple machines for the sake of efficiency. The fit of each takes 20 seconds to converge with 70 epochs. The triplane obtained from geometry reconstruction task and used for aerodynamics tasks is shown in Figure 1.
@@
Figure 1: Visulization of the raw triplane features of F_D_WM_WW_1282 (fastback with detailed underbody, with wheels, and with mirrors) used for surface field and 3D flow field predictions.
B.1.2 Triplane Resolution
@@
To evaluate the impact of triplane resolution on 3D geometry reconstruction, we compare different triplane resolutions and analyze their effects on reconstruction quality. Figure  2 illustrates the reconstructed 3D geometries at various triplane resolutions, while Table  2 provides quantitative metrics, including precision, recall, and F1-score, to assess reconstruction fidelity. Our results show that increasing triplane resolution improves reconstruction accuracy. At a lower resolution of $32\times 64\times 64$, the reconstructed geometry retains the overall shape but lacks fine surface details. As the resolution increases to $32\times 128\times 128$, the details become more refined, and the precision, recall, and F1-score improve. At $32\times 256\times 256$, the reconstruction achieves near-perfect fidelity, closely matching the original mesh, with an F1-score of 0.9936.
Figure 2: 3D Geometry reconstructed by triplanes with different dimensions.
@@
We compared the performance of model trained by triplanes with different dimensions and noticed that higher triplane resolutions come with increased memory and computational costs. As summarized in Table  3, A resolution of $32\times 128\times 128$ offers a good balance between reconstruction accuracy and efficiency. These results highlight the trade-off between resolution and computational efficiency, which should be carefully considered when selecting the appropriate triplane representation for downstream tasks.
@@
Table 2: Impact of triplane resolution on geometry reconstruction quality. Higher triplane resolutions lead to improved precision, recall, and F1 score, indicating more accurate and detailed geometry representation.
@@
Table 3: Performance of the Triplane network using different triplane resolutions on surface pressure field task on DrivAerNet dataset.
@@
B.2 Model Performance
The scalability of TripNet comes from the query-based prediction. All the baseline models need to build a complete graph to represent the geometry. However, the resolution and point number is limited by the GPU memory when it comes to meshes with millions of nodes in our experiment. TripNet is more flexible because all the geometry is represented by fixed dimensions of triplane feature maps. When the simulation resolution becomes higher, we use bilinear sampling to query features, which could still guarantee the accuracy while avoiding memory issues because the geometry representation cost is independent of mesh and simulation resolution.
Table 4: Scalability analysis on TripNet, FigConvNet, and Transolver.